Shift Summary for Owl, 8/26/99

Yee Bob Hsiung (hsiung@fnal.gov)
Thu, 26 Aug 1999 07:04:29 -0500 (CDT)

Shift Summary for Owl Shift, 08/26/99

On Shift: Ed Blucher, Bob Hsiung

Hours of beam available at all intensities: 3.5
Hours of beam available at nominal intensity: 3.
Hours of E832 data (special runs not included): 3.5
Hours of special runs (laser scans, muon runs...): 0.75

Accounting of lost time: 1 kfrend, missing spills, LINAC RF trip
and scheduled shutdown starts at 4am.

========================================================================
Runs ended this shift:
Run 14324 - BEAM_832 (16.8M events, 24.4K Vac00)
Run 14325 - Laser_scan

Tapes taken to Vault: QK*330-332, QKB115, QKN029
========================================================================
Accesses during shift: once to check TRD

========================================================================
Overview of shift:
3.5hrs E832 datataking (SEM=6.8E12) with some spills at 3E12.

Cleaned all tape drives except ktev3 backup drive.
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

14324 283 0:40 MCR tuning beam extraction attempts to smooth out
spill structures which give high DC1 current
from time-to-time. Intensity drop to~3E12.
292 0:52 Back to full intensity~6.6E12.
DC1 current is now ~100 uA.
1:00 MI instability, intensity vary.

333 1:46 ktevana error==>kfrend

While updating runlog summary entries for previous
run 14322 and 14323, we discovered that
yield 14322 gives 10734 total # spills?!?!
which corresponds to 238 hrs 32 mins or
9 days 22 hrs and 32 mins a real looooooooooong
owl shift Ed and I took yesterday!!

408 3:27 TeV abort.
3:30 abort again. LINAC RF tripped.
No more beam since.
14324 4:07 Stop run.Closed beam stops.12hrs Shutdown starts.

4:15 Install new CsI constants.
14325 4:17 Long Laser Scan.

4:45 Ramp down NM4AN with AN832NEGTOPOS.

14325 5:01 Xcellent. Unload QK*330-332, QKB115 + QKN029

5:30 Cleaned all tape drives except ktev3 backup.
Last cleaning was 7/30/99.
5:35 Ed and AG make access to check TRD.
========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

===>from Sasha G. (AG)
All 9 TRDs have been filled with Xe/CO2 and TRD gas recirculation
system has been turned ON.
SDAQ TRD-gas button goes to GREEN now.
===> DO NOT IGNORE TRD-gas RED button anymore <====

NM4AN flipped polarity to POS, ready for muon run.